OpenStudy (anonymous):

For n ∈ {0} ∪N, and k ∈ {0} ∪N, if k ≤ n, define C(n,k) = n!(k!)−1([n − k]!)−1. For n ∈ {0} ∪N, and k ∈ {0} ∪N, if k > n, define C(n,k) = 0. Prove that for any m,n ∈ {0} ∪N m+n sum k=m C(k,m) = C(m + n + 1,m + 1).

geerky42 (geerky42):

\(\displaystyle \sum_{k=m}^{m+n}C(k,m)=C(m + n + 1,m + 1)\) ?